OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10

OpenStudy (texaschic101):

we want to get the absolute value sign, and everything in it, on one side of the equal sign, and everything else on the other side. So we will first divide both sides by -5....giving us : | x + 1 | = -2 Now there will be 2 equations you will have to solve.. (x + 1) = -2 and -(x + 1) = -2 solve those for x and you will have your 2 solutions

OpenStudy (anonymous):

A. x = 0 B. x = −3 and x = 1 C. x = −1 and x = 3 D. No solution

OpenStudy (anonymous):

b

OpenStudy (anonymous):

the - is the negative sign

OpenStudy (anonymous):

thx @countryboy

OpenStudy (xapproachesinfinity):

incorrect there is no solution to that the absolute values is always positive therefore |x|=-a makes no sense then way no solutions